SHOW ALL THE WORK

2km 2km. 4.5 5km 1. Find the total resistance of the circuit. 2. Find the current through the 5 koresistor.
Answer #1

1. 2 kohm and 2 kohm are in series

R1 = (2+2) = 4 kohm

Now, 2 kohm and R1 are in parallel

(1/R2) = (1/2) + (1/4) = 3/4

R2 = 4/3 = 1.34 kohm

Now, R2 and 8 kohm are in series

R3 = R2 + 8 = 1.34+8 = 9.34 kohm

Now, R3 and 7 kohm are in parallel

1/R4 = (1/7)+(1/9.34)

R4 = 4.17 kohm

Now, R4 and 5 kohm in series

Req = 4.17+5 = 9.17 kohm

total resistance = 9.17 kohm

2. Current through 5 kohm,

V = IR

I = 4.5V/5 kohm = 9*10^-4 ampere